A die is rolled ten times find no. of ways so that outcomes always contain 1,2,3


 

Grade:11

1 Answers

Sher Mohammad IIT Delhi
9 years ago
probablity of getting of 1,2,3 =3/6=1/2
for 10 times =(1/2)^10


sher mohammad
askiitians
b.tech, iit delhi